OpenStudy (ageta):

An angle measures 128° less than the measure of a supplementary angle. What is the measure of each angle?

OpenStudy (ageta):

@CaseyCarns

OpenStudy (ageta):

@Catlover5925

OpenStudy (ageta):

@AnswerMyQuestions

OpenStudy (ageta):

@Directrix

Directrix (directrix):

Let x be one angle. Then, its supplement is 180 -x because supplementary angles sum to 180.

OpenStudy (ageta):

just give me the answer

Directrix (directrix):

An angle measures 128° less than the measure of a supplementary angle So, x = (180 - x) - 128

Directrix (directrix):

x = 180 - x - 128 2x = 52 x = 26 To get the second angle, compute 180 - 26 = ?

Directrix (directrix):

@ageta

OpenStudy (ageta):

tnxs